OpenStudy (anonymous):

Determine the equation of the line tangent to the graph of the following function at x = 0. g(x) = e^x(−4 + 3x + 3x^2)

OpenStudy (bahrom7893):

okay tangent line is the first derivative

OpenStudy (bahrom7893):

g(x) = e^x(−4 + 3x + 3x^2) g'(x) = (3+6x) * e^x(−4 + 3x + 3x^2)

OpenStudy (bahrom7893):

woops sorry the derivative was wrong..

OpenStudy (bahrom7893):

didn't see the x, btw is it (e^x)(−4 + 3x + 3x^2)?

OpenStudy (bahrom7893):

i mean is e^x separate?

OpenStudy (anonymous):

product rule?

OpenStudy (bahrom7893):

well yeah, but he doesn't specify... People use the freakin parenthesis, we have to guess what the heck was the original problem... Is (−4 + 3x + 3x^2) in the power too?

OpenStudy (bahrom7893):

i mean we are all busy students and u guys can't even make our lives easier?? Please, use a calculator notation, we will understand it, sorry if I sounded to harsh earlier, I was mad at someone else.. =)

OpenStudy (anonymous):

I think the equation of the tangent line is Y=-x-4

OpenStudy (anonymous):

first input the value of x as g(0) to fine the value of y. Then find the derivative using the product rule. factor out the e^x, and collect like terms. Then input the value of x into the derivative to find the slope. g'(x)=m. then use point slope formula.

OpenStudy (anonymous):

dear hix212, thank you for the help it was -x-4